OpenStudy (anonymous):

Help with differentiating f(x) = x^a - x^2a a is a positive integer I need to find out for what values of X the slope increases and decreases

OpenStudy (anonymous):

As in where the function f(x) is increasing and decreasing.

OpenStudy (anonymous):

you can call it f(x) = y = x^a - x^2a

OpenStudy (anonymous):

I find myself stuck around here after differentiating: \[a(x ^{a-1)}-2x ^{2(a-1)})\]

sam (.sam.):

\[a(x^{a-1}-2x^{2a-1})\]

OpenStudy (anonymous):

Yes of course, but I dont understand how to proceed to find out where the slope of the function is increasing/decreasing (and possibly a maximum and minimum)

sam (.sam.):

You can find the roots of it, then use values less than and more than the actual value to figure out whether it gives positive or negative result

sam (.sam.):

Here are the roots, \[\large x=( \frac{1}{2})^{\frac{1}{a}}\] \[\large x=0\] Use values larger or less than the roots to figure out whether its increasing or decreasing.